# Shadow of black holes with consistent thermodynamics

### v1

@@ VII Conclusions
In this paper, we have considered the minimal extension of the Kerr geometry that preserves the thermodynamic consistency, i.e., the line element (1) with a running Newton coupling $G(z)$. This minimal extension respects the circularity of the spacetime, which is a required spacetime symmetry in order to have a well-defined black hole temperature, and also ensures the Liouville integrability of geodesic dynamics. In particular, such an extension can be motivated by the approach of the asymptotic safety based on the FRG, in which the usual coupling constants in classical theory become scale-dependent. Modifications of the metric are also considered for the singularity resolution in black holes. The consistency of black hole thermodynamics further restricts the form of the running Newton coupling, i.e., Eq. (15). By considering a proper parameterization of the Newton coupling $G(z)$, we have investigated the shadow critical curves, that is, the shadow images contributed by infinitely lensed light rays, cast by this class of black hole models. The main aim of this study is to look for possible implications in terms of shadows if the black hole satisfies the thermodynamic consistency condition.
@@
Despite the minimal extension, the parameterization of the Newton coupling under consideration gives rise to a huge parameter space in our model. Under the variation of the coupling parameters in $G(z)$, as well as the spin of the black hole, the apparent size and the shape of the shadow critical curves change. This allows us to place constraints on the parameter space using the shadow size measurements available for the Sgr A* images (Figs. 3, 7, 9, and 11). In particular, the constraining power using supermassive black hole observations becomes extremely weak for the model in which the black hole singularity can be resolved by some putative quantum effects (Subsec. V.3). In addition, we find that the distortion in the shape of shadows, which is quantified by the circularity deviation $\Delta C$, is controlled by not only the spin $a_{*}$ and the inclination angle $\theta_{0}$, but also the coupling parameters in $G(z)$. In particular, for the parameter space we have chosen, the behavior of circularity deviation $\Delta C$ can be qualitatively captured by the extremality deviation parameter $\epsilon$ (Figs. 4, 8, 10, and 12), which is a quantity defined solely on the event horizon (Eq. (12)). The capability of $\epsilon$ to capture the behavior of the circularity deviation $\Delta C$ of the shadow critical curves in such a huge parameter space in our model is totally nontrivial. Moreover, it suggests the possibility of extracting geometrical information of the horizon scale from the shape of the critical curves when the black hole satisfies thermodynamic consistency.
@@
One particularly crucial result in this paper is the possible lower bound of the apparent shadow size in this model, as we discussed in Subsec. V.5. In our model, the size of shadow critical curves shrinks as the black hole spins up, or when one decreases the Newton coupling $G(z)$. For a given $a_{*}$, the shadow contour reaches its minimum size as the black hole approaches the extremal configuration, beyond which either the horizon disappears, or the shadow critical curves cannot maintain the D-shaped structure. We have explored as much as possible the parameter space of our model and all inclinations $\theta_{0}$, and have identified a lower bound of the fractional diameter deviation $\delta_{\textrm{min}}\gtrsim-0.21$ for black holes with $|a_{*}|\leq 1$ (Fig. 13). Such a universal lower bound of $\delta$ could be used as an observational indicator to test the consistency of black hole thermodynamics. More explicitly, if there exist black holes whose D-shaped shadows have $\delta$ violating this lower bound, such black holes must violate at least one of the assumptions made in our paper, which include:
  • The setting of the minimal extension, i.e., the line element Eq. (1) with a running $G(G_{0}/(r^{2}+a^{2}),Mr)$.
  • The parameterization for $G(G_{0}/(r^{2}+a^{2}),Mr)$.
@@
One possible extension of this work is to go beyond the minimal setting in Eq. (1) by considering, for instance, the most general Liouville integrable spacetime Benenti:1979erw ; Papadopoulos:2018nvd . One can even relax the assumption of Liouville integrability and consider a more general spacetime parameterization Konoplya:2016jvv . It is highly desirable to identify the conditions for thermodynamic consistency in these general spacetime settings, and to look for possible observational implications. We will leave these interesting issues to future work.
Acknowledgement

### v2

@@ VII Conclusions
In this paper, we have considered the minimal extension of the Kerr geometry that preserves the thermodynamic consistency, i.e., the line element (II) with a running Newton coupling $G(z)$. This minimal extension respects the circularity of the spacetime, which is a required spacetime symmetry in order to have a well-defined black hole temperature, and also ensures the Liouville integrability of geodesic dynamics. In particular, such an extension can be motivated by the approach of the asymptotic safety based on the FRG, in which the usual coupling constants in classical theory become scale dependent. Modifications of the metric are also considered for the singularity resolution in black holes. The consistency of black hole thermodynamics further restricts the form of the running Newton coupling, i.e., Eq. (15). By considering a proper parametrization of the Newton coupling $G(z)$, we have investigated the shadow critical curves, that is, the shadow images contributed by infinitely lensed light rays, cast by this class of black hole models. The main aim of this study is to look for possible implications in terms of shadows if the black hole satisfies the thermodynamic consistency condition.
@@
Despite the minimal extension, the parametrization of the Newton coupling under consideration gives rise to a huge parameter space in our model. Under the variation of the coupling parameters in $G(z)$, as well as the spin of the black hole, the apparent size and the shape of the shadow critical curves change. This allows us to place constraints on the parameter space using the shadow size measurements available for the Sgr A* images (Figs. 3, 7, 9, and 11). In particular, the constraining power using supermassive black hole observations becomes extremely weak for the model in which the black hole singularity can be resolved by some putative quantum effects (Sec. V.3). In addition, we find that the distortion in the shape of shadows, which is quantified by the circularity deviation $\Delta C$, is controlled by not only the spin $a_{*}$ and the inclination angle $\theta_{0}$ but also the coupling parameters in $G(z)$. In particular, for the parameter space we have chosen, the behavior of circularity deviation $\Delta C$ can be qualitatively captured by the extremality deviation parameter $\epsilon$ (Figs. 4, 8, 10, and 12), which is a quantity defined solely on the event horizon (Eq. (12)). The capability of $\epsilon$ to capture the behavior of the circularity deviation $\Delta C$ of the shadow critical curves in such a huge parameter space in our model is totally nontrivial. Moreover, it suggests the possibility of extracting geometrical information of the horizon scale from the shape of the critical curves when the black hole satisfies thermodynamic consistency.
@@
One particularly crucial result in this paper is the possible lower bound of the apparent shadow size in this model, as we discussed in Sec. V.5. In our model, the size of shadow critical curves shrinks as the black hole spins up, or when one decreases the Newton coupling $G(z)$. For a given $a_{*}$, the shadow contour reaches its minimum size as the black hole approaches the extremal configuration, beyond which either the horizon disappears, or the shadow critical curves cannot maintain the D-shaped structure. We have explored as much as possible the parameter space of our model and all inclinations $\theta_{0}$ and have identified a lower bound of the fractional diameter deviation $\delta_{\textrm{min}}\gtrsim-0.21$ for black holes with $|a_{*}|\leq 1$ (Fig. 13). Such a universal lower bound of $\delta$ could be used as an observational indicator to test the consistency of black hole thermodynamics. More explicitly, if there exist black holes whose D-shaped shadows have $\delta$ violating this lower bound, such black holes must violate at least one of the assumptions made in our paper, which include:
  • The setting of the minimal extension, i.e., the line element Eq. (II) with a running $G(G_{0}/(r^{2}+a^{2}),Mr)$.
  • The parametrization for $G(G_{0}/(r^{2}+a^{2}),Mr)$.
@@
We would like to emphasize that although the spacetime metric (II) with a running Newton coupling preserves the Liouville integrability of geodesic dynamics, which implies that the model is a subclass of the general metric parametrizations considered in Refs. Johannsen:2013szh ; Papadopoulos:2018nvd , the restriction on the form of the running Newton coupling required by the consistency of thermodynamics significantly reduces the arbitrariness of the metric parametrizations. Therefore, although the parameter space of our model is still large, the model cannot generate arbitrary shadows, and that is why we have been able to identify common shadow features (e.g., the behavior of $\epsilon$ and the universal lower bound of $\delta$) shared in such a huge parameter space. These shadow features may not exist anymore when going beyond the subclass our model belongs to, where the consistency of thermodynamics may be broken.
One possible extension of this work is to go beyond the minimal setting in Eq. (II) by considering, for instance, the most general Liouville integrable spacetime Benenti:1979erw ; Papadopoulos:2018nvd . One can even relax the assumption of Liouville integrability and consider a more general spacetime parametrization Konoplya:2016jvv . It is highly desirable to identify the conditions for thermodynamic consistency in these general spacetime settings, and to look for possible observational implications. We will leave these interesting issues to future work.
Acknowledgements